OpenStudy (anonymous):

In the below system, solve for y in the first equation. x − 3y = −6 2x − 7y = 10

OpenStudy (lieutenantgeneral):

y = 3 n and/or n + 2

OpenStudy (anonymous):

Choices are 1/3x + 2 -1/3x + 6 −x + 2 -1/3x + 2

OpenStudy (lieutenantgeneral):

-1/3x + 6

OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x − 3y = −3 2x − 6y = −6

OpenStudy (lieutenantgeneral):

5x − 3y = −3 = y = 1/3(5 x + 3) 2x − 6y = −6 = y = x + 3/3

OpenStudy (anonymous):

answer choices −1 0 1 2

OpenStudy (lieutenantgeneral):

1

OpenStudy (lieutenantgeneral):

i meant -1

OpenStudy (anonymous):

Can I message you